OpenStudy (anonymous):

Jorge’s bowling score was 230 last week, but only 180 this week. What was the percent decrease in Jorge’s score to the nearest percent? A. 128% B. 78% C. 28% D. 22%

OpenStudy (anonymous):

I think d, can someone check, i will give medal

OpenStudy (dan815):

Hi!

OpenStudy (anonymous):

hello

OpenStudy (dan815):

you had 230 and you lost some perfect of 230 to give 180: lets write that in math

OpenStudy (anonymous):

kk

OpenStudy (dan815):

230 - x% of 230 = 180 230- x/100 * 230 = 180 230(1-x/100)=180 lets solve for x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got 500/23

OpenStudy (calculusxy):

Here is a formula, personally my favorite: \[\frac{ greater - lesser }{ original } \times 100\]

OpenStudy (dan815):

230*(1-x/100)=180 230*R=180 R=180/230 1-x/100=180/230 =230/230 - 50/230 1-x/100= 1-50/230 we al most have it looking identical can you do the rest?

OpenStudy (dan815):

what must x be?

OpenStudy (dan815):

yep 500/23 is right :)

OpenStudy (anonymous):

do i have to change that into a percent?

OpenStudy (anonymous):

2173.913%

OpenStudy (anonymous):

i personaly think it is d, is that it?

OpenStudy (dan815):

500/23 = ?

OpenStudy (calculusxy):

@Houdini_Dragon I think you're right.

OpenStudy (anonymous):

alrighty, thanks so much, it is d thanks dan and cal!

OpenStudy (dan815):

you're welcome :)